OpenStudy (anonymous):

Solve for x : |2x + 4| = 16

OpenStudy (mathteacher1729):

Have you tried graphing this?

OpenStudy (anonymous):

no i really dont know what to do

OpenStudy (anonymous):

\[|2x+4|=16\] \[2x+4=16\] \[2x+4=-16\]

OpenStudy (anonymous):

solve each separately \[2x+4=16\] \[2x=12\] \[x=6\] or \[2x+4=-16\] \[2x=-20\] \[x=-10\]

OpenStudy (mathteacher1729):

To graph: 1 go to http://www.wolframalpha.com 2 type "plot y =|2x+4| and y = 16"